OpenStudy (anonymous):

differentiate e^(x/2)

OpenStudy (anonymous):

e^(x/2) * 0.5

OpenStudy (anonymous):

differentiate of e^(x/2) = the same function * differentiate of the power; e^(x/2) * 0.5

OpenStudy (michele_laino):

if y=e^(x/2), then the first derivative y', is equal to: \[y'=e ^{(x/2)}*\frac{ 1 }{ 2 }\] it is necessary to apply the chain rule of derivatives